[Brahms-l] Shift report 20040314 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sun Mar 14 2004 - 15:56:53 EST
------------------------------------------------------------
Date: Sun Mar 14 15:56:51 2004, Shift: 08:00-16:00
Supervisor: JHL, with: 
Log book#15 pages 87 - 87

Shift Summary:
------------------------------------------------------------
Beam:
- 1 full store for the shift
- 4.5 hrs (56 bunches) with low intensity 
  (ZDC ~ 2600/sec -> 600/sec)

Data:
- Runs 11007 - 11010
- FS 8deg 1/1A: 128K trigger 2
- MRS 90deg 12B: 166K trigger 3
- Min-bias Run 11013: 370K trigger 4

Problems: 
- TRMRS TDCs are back 
  (see the e-log for the details) 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 14 15:57:09 2004

This archive was generated by hypermail 2.1.8 : Sun Mar 14 2004 - 15:57:25 EST